OpenStudy (rainbow_dash):

Find the sum of the given arithmetic series. 24 + 48 + 72 + 96 + ... + 480

OpenStudy (anonymous):

Do you remember what the formula is?

OpenStudy (rainbow_dash):

Well my class is based on pretests and usually I forget what I do once i pass each test..but i have a_n=a_1+(n-1)d and for the sum S_n=n(a_1+a_n)/2

OpenStudy (anonymous):

Yeah so just start with finding n, using the first expression you gave me

OpenStudy (anonymous):

Then sub that value of n into the second equation c:

OpenStudy (rainbow_dash):

Okay thanks :)

OpenStudy (anonymous):

Hope this helped c:

OpenStudy (anonymous):

Oh sorry had a derp moment.. Remember how a_n can be (a_4) - (a_3) = (a_2) Or (a_3) - (a_2) = (a_1) You can use it to help find n